OpenStudy (anonymous):

lim(p→3)⁡(4x^2+5x-9)

OpenStudy (amistre64):

there is content missing

OpenStudy (amistre64):

lol, or its a convergence question ...

OpenStudy (anonymous):

the limit of p as it approaches 3

OpenStudy (anonymous):

this is correct

OpenStudy (amistre64):

there is no p in your rule to assess it

OpenStudy (amistre64):

lim(p→3)⁡(4x^2+5x-9) ; the value of p has no effect on this rule

OpenStudy (anonymous):

hmm your right lol this must of been a mistake

OpenStudy (anonymous):

lim(x→4)⁡(x^2-16)/(x-4)

OpenStudy (anonymous):

how bout this one

OpenStudy (amistre64):

thats more cogent :)

OpenStudy (amistre64):

lets test something out, what is the value of the top and bottom when x = 4?

OpenStudy (anonymous):

@Toadil do you know what a difference of squares is?

OpenStudy (anonymous):

lim(x→4)⁡(4^2-16)/(0) undefined

OpenStudy (amistre64):

4^2 - 16 = 0 4 - 4 = 0 since both equate to zero, this tell us that they have a common factor get them into their factored form to cancel out the like factor and reassess it

OpenStudy (amistre64):

since (x-4) is not factorable, then the top must have a factor of (x-4) hidden in it

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

are you able to factor the top? oldrin suggested its a difference of squares if that rings a bell; if not, then you would have to rely upon a more basic method

OpenStudy (amistre64):

or some fancier new ones lol

OpenStudy (anonymous):

(x^2-16) = 2(x - 8)

OpenStudy (amistre64):

not quite, x^2 is not the same as 2x so that is not a proper application of factoring

OpenStudy (amistre64):

we know that (x-4) is a factor (x-4)(x+n) = x^2-16 , distribute the left side (x-4)x+(x-4)n = x^2-16 , distribute the left side again xx-4x+nx-4n = x^2-16 , add 0x to the right side for comparison x^2 +(n-4)x -4n = x^2 + 0x -16 , and compare parts x^2 +(n-4)x -4n x^2 + 0 x -16 n-4 = 0 and -4n = 16 solve for n

OpenStudy (amistre64):

if we go back to grade school and just divide it out ... x + 4 <-- our other factor ------------- x-4 | x^2 - 16 -(x^2 -4x) ---------- 4x - 16 -(4x - 16) --------- 0 + 0 <-- nothing remains therefore; (x-4)(x+4) = x^2 - 16

OpenStudy (amistre64):

so, what does this do for us? \[\frac{x^2-16}{x-4}=\frac{\cancel{(x-4)}(x+4)}{\cancel{x-4}}=x+4\] \[\lim_{x \to4}~(x+4)=(4+4)\]

OpenStudy (anonymous):

As x approaches 4 x equals 4

OpenStudy (amistre64):

as x approached 4, the function approaches 8

OpenStudy (anonymous):

oh duh lol I was think x = 4

OpenStudy (anonymous):

so the functions would equal 8 lol

OpenStudy (anonymous):

4 + 4

OpenStudy (amistre64):

graphically,

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

the value of the function does not exist at x=4, but the limit exists as x approaches 4

OpenStudy (anonymous):

Okay thank you very much for your help.

OpenStudy (amistre64):

youre welcome